OpenStudy (anonymous):

Need help in geometry! DF bisects angle EDG. Find FG. The diamgram is not to scale.

OpenStudy (anonymous):

EF = n + 5 FG = 2n-6

OpenStudy (anonymous):

Please show me the steps as to how you solved it, thx

OpenStudy (decarr432):

@mathmale

OpenStudy (anonymous):

Do you have any idea? I know that FG = FE

OpenStudy (anonymous):

im just not sure how to solve the equation

OpenStudy (loser66):

Let them equal and solve for n

OpenStudy (anonymous):

is the answer n = 11?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Yup

OpenStudy (anonymous):

Thank you!!

OpenStudy (loser66):

but n =11 is not the final answer, right? because EF = n + 5 , hence EF =?

OpenStudy (anonymous):

16?

OpenStudy (loser66):

yup, and your question: FG =?

OpenStudy (anonymous):

16 Thank you again:)

OpenStudy (loser66):

ok